Consider the titration of 40.0 mL of 0.0600 M C2H5NH2 (a weak base; Kb = 0.000640) with 0.100 M HCl. Calculate the pH after the following volumes of titrant have been added:

(a) 0.0 mL pH =

(b) 6.0 mL pH =

(c) 12.0 mL pH =

(d) 18.0 mL pH =

(e) 24.0 mL pH =

(f) 26.4 mL pH =
Answer #1

And given: volume Molarity of (245NH2 = 40mL of C245 MH2 = 0.06oom =0.0 Goomall Molarity of HCl = 0.100M = 0 loomalle (a) volSolving for a a= 0.00588503 Thus, (on] - 0-00588503 pok= -log Core] = -log (0.00588503) = 2.230 25132 pH = 14-pon =14 – 2230= 0.0024 mol to the ² C2H5NH3 reaction + on- Constructing ICE table C2H5NH2 +420 Initial 0.0024 (moles) Change - 0.0006 Equil-0.0018 468103 -0.0391 304348 Poll= Pkb Applying Henderson- Hasselbalch equation: Pol= pkb + log [C2kg MH37] [.45 мм) I K6= 0Thus, Moles ginclo / Ricles of Chanta) This is midpoint Thus, pola pkb = 3.1938 2003 PH= 14 - pon = 14- 3.1938 2003 = 10.8061(C215042] = (0.0024 -0.0018) mol 58x1034 = 0.0006 M20.0103 44 8276 M 58x1ő3 [Ca H5NH3+] = 0.0018 Mc 0.0310 344828M 58x103 AppThis is equivalence point Constructing ICE table C2H5NH2 +420 Chittal for the reaction: Ą C2H5rikat ton (moles) 0.0024 ChangeКь - t-00xo aliooxic14 о: ooo 640 11. 5625 xia Ка: Сери, чи,) Cugo) Со Mқми,t) 2 (-5325х1 = О-о за5- - (3 — Б. 85a315xo 20(m volume of Hal = 26.4mL = 26.4x103 Moles of real = 26.4x10²4 x o. looma) = 0.00264 mol Excers Moles of rel= Moles of real-